OpenStudy (anonymous):

Find the solution of the exponential equation, correct to four decimal places. 5^x=4^(x+2)

OpenStudy (anonymous):

\(\large 5^x=4^{x+2} \) \(\large 5^x=4^x \cdot 4^2 \) \(\large \frac{5^x}{4^x}= 4^2 \) \(\large (\frac{5}{4})^x= 2^4 \) \(\large x\cdot ln(\frac{5}{4})=4ln2 \) \(\large x\cdot (ln5-ln4)=4ln2 \) \(\large x\cdot (ln5-2ln2)=4ln2 \) \(\large x=\frac{4ln2}{ln5-2ln2} \)

OpenStudy (anonymous):

now just use your calculator....

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!!! I WOULD HAVE NEVER THOUGHT OF DOING THAT!!

OpenStudy (anonymous):

yw... i don't have a calculator but wolfram will do.... http://www.wolframalpha.com/input/?i=solve+5%5Ex%3D4%5E%28x%2B2%29